----------------------------------------------------------------------------------------------------------------------------------- Attempt all question
Q1: During the last few years, Harry Davis Industries has been too constrained by the high cost of capital to make many capital investments. Recently, though, capital costs have been declining, and the company has decided to look seriously at a major expansion program that has been proposed by the marketing department. Assume that you are an assistant to Leigh Jones, the financial vice president. Your first task is to estimate Harry Davis’s cost of capital. Jones has provided you with the following data, which she believes may be relevant to your task: (1) The firm’s tax rate is 40%.
(2) The current price of Harry Davis’s 12% coupon, annual payment, bonds with 15 years remaining to maturity is $1,153.72. Harry Davis does not use short-term interest-bearing debt on a permanent basis. New bonds would be privately placed with no flotation cost.
(3) The current price of the firm’s 10%, $100 par value, annual dividend, perpetual preferred stock is $116.95. Harry Davis would incur flotation costs equal to 5% of the proceeds on a new issue.
(4) Harry Davis’s common stock is currently selling at $50 per share. Its last dividend was $4.19, and dividends are expected to grow at a constant rate of 5% in the foreseeable future. Harry Davis’s beta is 1.2, the yield on T-bonds is 7%, and the market risk premium is estimated to be 6%. For the bond yield- plus- risk-premium approach, the firm uses a 4 percentage point risk premium.
(5) Harry Davis’s target capital structure is 30% long-term debt, 10% preferred stock, and 60% common equity.
To structure the task somewhat, Jones has asked you to answer the following questions.
What is Harry Davis’s weighted average cost of capital (WACC)?
 Why is cost of capital calculation important for a company and how can it help the investors?
Answer #1
1..After-tax cost of annual 12% coupon bonds, kd
Using the formula to find the price of the bond,
ie. Price=Present value of future coupons+PV of face value to be received at maturity
ie. Price=(Annual Coupon*(1-(1+r)^-n)/r)+(FV/(1+r)^n
Plugging in the given values,
1153.72=((1000*12%)*(1-(1+r)^-15)/r)+(1000/(1+r)^15)
Solving for r, we get the before-tax cost of the bond as,
9.98116%
Now, the after-tax cost=Before-tax cost*(1-Tax rate)
ie. 9.98116%*(1-40%)=
5.99%
Cost of preferred stock, kps
kps=$ Annual dividend/Net proceeds of issue
where,Net proceeds of issue=Market price-Flotation costs
ie. (100*10%)/(116.95*(1-5%))=
9.00%
Cost of commmon stock,ke
As per Gordon's constant growth model of dividend cash flows
ke=(D1/P0)+g
where, D1=the next dividend, ie. D0*(1+g),ie.4.19*(1+5%)= $ 4.3995
P0= $ 50 & g= 5%
So,ke=(4.3995/50)+5%=
13.80%
Cost of equity , ke,Using the CAPM approach,
ke=Risk-free rate+(Beta*Market risk premium)
ie. 7%+(1.2*6%)=
14.20%
Cost of Equity, ke, using bond yield- plus- risk-premium approach
ke= Bond Yield + 4%
we will use after-tax bond yield , as every $ for equity is after-tax only
so, 5.99%+4%=
9.99%
Now, averaging all the above 3 costs of equity,
we get the average cost of equity, ke , as
(13.80%+14.20%+9.99%)/3=
12.66%

Harry Davis’s weighted average cost of capital (WACC)

WACC=(Wt.d*kd)+(Wt.ps*kps)+(wt.e*ke)
(30%*5.99%)+(10%*9%)+(60%*12.66%)=
10.29%
(ANSWER)
2.Cost of capital calculation helps the investor to know the cost of his funds employed in a project or any specific venture , so that he can compare the net return % from this project to the cost of capital.
If he can calculate & have the cost of his sources of finance , at hand, it will help him to monitor the returns on his investments .That way ,he can avoid ungainful investments at the start itself.
Knowing the cost of capital of his potential investment companies will help him to evaluate the company's cash flows & arrive at its current value.
It is important for an investor to analyse how far or near is the cost of capital to his required rate of return , in any investment opportunity, he undertakes.